# FRAMES: Boosting LLMs with A Four-Quadrant Multi-Stage Pretraining Strategy

### v1

@@ Abstract
Large language models (LLMs) have significantly advanced human language understanding and generation, with pretraining data quality and organization being crucial to their performance. Multi-stage pretraining is a promising approach, but existing methods often lack quantitative criteria for data partitioning and instead rely on intuitive heuristics. In this paper, we propose the novel Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES), guided by the established principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. This principle is grounded in two key findings: first, training on high Perplexity (PPL) data followed by low PPL data, and second, training on low PPL difference (PD) data followed by high PD data, both causing the loss to drop significantly twice and performance enhancements. By partitioning data into four quadrants and strategically organizing them, FRAMES achieves a remarkable 16.8% average improvement over random sampling across MMLU and CMMLU, effectively boosting LLM performance.
FRAMES: Boosting LLMs with
@@ 1 Introduction
Figure 1: Few-shot performance for 3B models shows FRAMES achieves 15.3% improvement over Random across MMLU, and 18.2% on CMMLU.
In this paper, we propose a novel Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES) to boost LLMs’ performance, which is guided by the principle of organizing the pretraining process into four stages to achieve significant loss drops four times. The principle is derived from two key findings based on quantifiable data metrics.
@@
Based on these two major findings, we establish the principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. Specifically, we partition the data into four quadrants based on PPL and PD: Quadrant 1 ($Q_{1}$) with low PD and low PPL, Quadrant 2 ($Q_{2}$) with high PD and low PPL, Quadrant 3 ($Q_{3}$) with low PD and high PPL, and Quadrant 4 ($Q_{4}$) with high PD and high PPL. Through further analysis and experiments, we determine that the optimal strategy is to reorganize the pretraining data in the $Q3\to Q4\to Q1\to Q2$ sequence. We also implement an inner-to-outer hierarchical gradual transition method to ensure smooth data transitions between stages. Practical evidence shows that FRAMES results in loss dropping significantly four times, enhances the models’ emergent abilities, and boosts their performance.
In summary, our contributions are as follows: (1) We identify two key findings: first, training on high PPL data followed by low PPL data, and second, training on low PD data followed by high PD data, both causing the loss to drop significantly twice and performance enhancements. (2) Guided by the principle of continuous loss reduction, we propose FRAMES, a novel strategy that partitions data into four quadrants and organizes the pretraining process into four stages, resulting in loss dropping significantly four times and boosting model performance. (3) Experiments demonstrate a 16.8% improvement over uniform sampling in average performance across MMLU and CMMLU.
@@ 2 Four Quadrant Multi-stage Pretraining Strategy
In this section, we first evaluate the effectiveness of PPL and PD as metrics for data organization in two-stage training (Sections 2.1 and 2.2). Based on these insights, we propose FRAMES with both metrics(Section 2.3), as shown in Figure 2.
@@ 2.1 Two-stage pretraining guided by PPL
(a) PPL training loss
(b) PD training loss
@@
Random 24.84 25.57
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.96 25.97
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.63 42.63
@@ 2.2 Two-stage pertaining guided by PD
where $PPL_{M_{w}}(x)$ and $PPL_{M_{s}}(x)$ are the perplexity values of the sample $x$ calculated using $M_{w}$ and $M_{s}$, respectively. A low PD indicates similar learning efficiency for both models, while a high PD suggests the sample is more challenging for the weak model.
@@
The distributions of PPL and PD on different domains are analyzed in Appendix B.5, PD exhibits a relatively consistent distribution across varied domains, following a normal distribution with a mean value of 0.3. Compared to PPL, PD offers a better advantage in maintaining data diversity throughout each stage of training.
@@
We train a 100M parameter RM and then calculate the PD for each sample using both the 100M and 1.3B RMs, referred to as PD(100M-1.3B). Using the median value of PD across all data, we divide the training dataset into two subsets with equal token counts: $\text{A}_{\text{PD}}^{\text{low}}$ and $\text{A}_{\text{PD}}^{\text{high}}$. We conduct a two-stage training process with $\text{A}_{\text{PD}}^{\text{low}}$ first, followed by $\text{A}_{\text{PD}}^{\text{high}}$, and compare the results to the random setting. We use the evaluation methods described in Section 2.1.
@@
Random 24.84 25.57
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.89 27.20
@@ Results
Figure 3(b) shows the training loss changes under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setup. The loss initially drops rapidly during the low PD phase, then stabilizes, and decreases further with high PD data, eventually falling below the Random model’s loss. It suggests that the first phase sets a beneficial optimization path for the second, helping avoid local optima. Table 2 shows the model’s accuracy under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. Notably, it exceeds the Random setting by 2.05% on MMLU and 1.63% on CMMLU, which validates PD as an effective metric. Further, this leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
@@ 2.3 Four-Quadrant Guided Training Strategy
Inspired by the two key findings about PPL and PD, we establish the principle of organizing pretraining data to achieve significant reductions in training loss. Based on this principle, we introduce a novel Four quadRAnt Multi-stage prEtraining Strategy (FRAMES), which uses PPL and PD to partition data and reorganize the data sequence to make the training loss drop significantly. The process of FRAMES is shown in Algorithm 1.
Specifically, we train two RMs on the target training set $D$, with the strong model $M_{s}$ having more parameters than the weak model $M_{w}$. Both models are trained on data from the same distribution and under identical settings. We compute PPL using $M_{s}$ and PD using both $M_{s}$ and $M_{w}$ for each data point in $D$. We use the two metrics to partition $\mathcal{D}$ into four quadrants, as shown in Figure 4(a). The data from four quadrants can be described as follows: Quadrant $\mathbf{Q_{1}}$ contains data that both $M_{s}$ and $M_{w}$ can fit well; Quadrant $\mathbf{Q_{2}}$ contains data that $M_{s}$ learns well, but $M_{w}$ struggles with. Data from this subset is challenging and requires a higher model capacity to understand them; Quadrant $\mathbf{Q_{3}}$ contains data poorly learned by both $M_{s}$ and $M_{w}$; Quadrant $\mathbf{Q_{4}}$ contains data poorly learned by $M_{s}$ and even worse by $M_{w}$.
@@
Figure 4: Pretraining losses of four quadrants and main experiments.
@@
To efficiently organize data from the four quadrants while adhering to constraints, we use a double-loop approach: first, dividing the training into two main phases based on a specific constraint, then further splitting each main phase into two sub-phases according to another constraint. The approach yields two distinct four-stage training strategies: the first strategy follows the sequence $Q_{3}\rightarrow Q_{4}\rightarrow Q_{1}\rightarrow Q_{2}$, while the second strategy adopts $Q_{3}\rightarrow Q_{1}\rightarrow Q_{4}\rightarrow Q_{2}$. Both strategies break the constraints between the second and third stages. However, experiments in Sections 2.1 and 2.2 show that the loss change from low to high PPL is much greater than from high to low PD. Thus, we prioritize the constraints in Equation (2) over that in Equation (3), making the first four-stage training strategy $\mathbf{Q_{3}\to Q_{4}\to Q_{1}\to Q_{2}}$ the better choice. Furthermore, we implement an inner-to-outer hierarchical gradual transition method to ensure smooth data transitions between stages.
Algorithm 1 FRAMES
1:Input: pretraining data $D$, all training steps $K$, smooth steps $n$
2:Output: trained model $\theta_{K}$
3:Initialize model parameters $\theta_{0}$
4:Train RMs on i.i.d. subset of $D$
5:Calculate PPL and PD for all samples in $D$ using RMs
6:Divide $D$ into 4 quadrants: $Q_{1}$, $Q_{2}$, $Q_{3}$, and $Q_{4}$, based on PPL and PD
7:Inner-layer mixing:
8:for $t=1$ to $\frac{|Q_{3}|+|Q_{4}|}{N}$ do
9:     Sample batch $B_{t}^{34}$ from $Q_{3}$ and $Q_{4}$
10:end for
11:$S_{34}\leftarrow\bigcup_{t=1}^{\frac{|Q_{3}|+|Q_{4}|}{N}}B_{t}^{34}$
12:for $t=1$ to $\frac{|Q_{1}|+|Q_{2}|}{N}$ do
13:     Sample batch $B_{t}^{12}$ from $Q_{1}$ and $Q_{2}$
14:end for
15:$S_{12}\leftarrow\bigcup_{t=1}^{\frac{|Q_{1}|+|Q_{2}|}{N}}B_{t}^{12}$
16:for $k=1$ to $K$ do
17:     Outer-layer mixing:
18:     Sample final batch $B_{k}$ from $S_{34}$ and $S_{12}$
19:     Update model parameters $\theta_{k}$ using batch $B_{k}$
20:end for
Formulation of FRAMES with Stage Transition
Given the data organization sequence $S=[Q_{3},Q_{4},Q_{1},Q_{2}]$, we use the S-shape function from PDPCZhang et al. (2025) to organize the data as a hierarchical two-source data mixing process:
$f(p)=\frac{1}{1+exp{(a(p-0.5))}}$ (4)
@@
where $p\in[0,1]$, represents the completion ratio of the pretraining process. $f(p)$ provides a proportion of data from the first source, and $a$ modulates the steepness of the curve. Unlike PDPC, we merely utilize the S-shape function for smoothing during stage transitions. Therefore, we use a larger $a=35$ instead of $a=10$ as in PDPC, to achieve a steeper function curve, as illustrated in Appendix B.4.
Formally, the inner-layer mixing is expressed as:
$B_{t}^{ij}=\{x\mid x\sim Q_{i}\}_{f(p)\cdot N}\cup\{x\mid x\sim Q_{j}\}_{(1-f(p))\cdot N}$ (5)
where N is the batch size. When $i=3$ and $j=4$, the mixed data from $Q_{3}$ and $Q_{4}$ at step $t$ is denoted as $B_{t}^{34}$; similarly, when $i=1$ and $j=2$, the mixed data from $Q_{1}$ and $Q_{2}$ at step $t$ is denoted as $B_{t}^{12}$. We take the union of the batches from the inner-layer data mixes to form the outer-layer mixed data sources $S_{34}$ and $S_{12}$, respectively:
$S_{34}=\bigcup_{t=1}^{\frac{|Q_{3}|+|Q_{4}|}{N}}B_{t}^{34},\quad S_{12}=\bigcup_{t=1}^{\frac{|Q_{1}|+|Q_{2}|}{N}}B_{t}^{12}$ (6)
Then the outer-layer mixing is then expressed as:
$B_{t}=\{x\mid x\sim S_{34}\}_{f(p)\cdot N}\cup\{x\mid x\sim S_{12}\}_{(1-f(p))\cdot N}$ (7)
@@
where $f(p)$ controls the proportions of each part during the inner and outer-layer mixing processes. Notably, FRAMES only organizes the given data without performing selection, which can serve as the final data preprocessing step before pretraining.
@@ 3 Experiments
Metric Order ARC-E ARC-C SciQ HellaSw. PIQA AVG.
- Random 56.5 23.6 85.8 34.2 67.3 53.5
PD $\downarrow$ 54.7$\downarrow$1.8 21.8$\downarrow$1.8 87.1$\uparrow$1.3 33.7$\downarrow$0.5 67.8$\uparrow$0.5 53.0$\downarrow$0.5
PD $\uparrow$ 56.1$\downarrow$0.4 21.3$\downarrow$2.3 86.2$\uparrow$0.4 34.4$\uparrow$0.2 67.6$\downarrow$0.2 53.1$\downarrow$0.4
PPL $\downarrow$ 45.5$\downarrow$11.0 20.6$\downarrow$3.0 71.2$\downarrow$14.6 30.3$\downarrow$3.9 63.7$\downarrow$3.6 46.3$\downarrow$7.2
PPL $\uparrow$ 47.8$\downarrow$8.7 17.9$\downarrow$5.7 72.7$\downarrow$13.1 29.1$\downarrow$5.1 62.4$\downarrow$4.9 46.0$\downarrow$7.5
Qu.Edu $\downarrow$ 57.2$\uparrow$0.7 26.4$\uparrow$2.8 79.4$\downarrow$6.4 31.0$\downarrow$3.2 62.7$\downarrow$4.6 51.3$\downarrow$2.2
Qu.Edu $\uparrow$ 56.8$\uparrow$0.3 26.0$\uparrow$2.4 82.1$\downarrow$3.7 32.0$\downarrow$2.2 64.9$\downarrow$2.4 52.4$\downarrow$1.1
PPL S.R. 56.1$\downarrow$0.4 24.1$\uparrow$0.5 87.8$\uparrow$2.0 33.9$\downarrow$0.3 67.4$\uparrow$0.1 53.9$\uparrow$0.4
@@
Qu.Edu S.R 54.7$\downarrow$1.8 22.9$\downarrow$0.7 86.2$\uparrow$0.4 33.6$\downarrow$0.6 67.4$\uparrow$0.1 53.0$\downarrow$0.5
Qu.Edu S. 54.5$\downarrow$2.0 22.8$\downarrow$0.8 87.8$\uparrow$2.0 34.0$\uparrow$0.2 67.9$\uparrow$0.6 53.4$\downarrow$0.1
PDPC S. 57.3$\uparrow$0.8 26.6$\uparrow$3.0 87.9$\uparrow$2.1 33.7$\downarrow$0.5 68.0$\uparrow$0.7 54.7$\uparrow$1.2
FRAMES - 62.9$\uparrow$6.4 26.5$\uparrow$2.9 90.5$\uparrow$4.7 38.4$\uparrow$4.2 69.6$\uparrow$2.3 57.6$\uparrow$4.1
Table 3: Downstream tasks results for different settings on 1.3B models. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: AVG. = Average.
Method MMLU CMMLU CEVAL BBH ARC-C PIQA AVG.
Random 27.7 27.5 27.2 27.9 33.7 76.0 36.7
PDPC 35.8$\uparrow$8.1 35.6$\uparrow$8.1 36.1$\uparrow$8.9 25.7$\downarrow$2.2 35.8$\uparrow$2.1 76.3$\uparrow$0.3 40.9$\uparrow$4.2
$Q_{3}\to Q_{1}\to Q_{4}\dashrightarrow Q_{2}$ 25.8$\downarrow$1.9 26.4$\downarrow$1.1 25.5$\downarrow$1.6 28.9$\uparrow$1.0 35.2$\uparrow$1.5 75.4$\downarrow$0.6 36.5$\downarrow$0.2
FRAMES 43.0$\uparrow$15.3 45.7$\uparrow$18.2 44.0$\uparrow$16.8 27.90.0 36.5$\uparrow$2.8 76.9$\uparrow$0.9 45.7$\uparrow$9.0
@@
(a) Accuracy on CEVAL.
(b) Accuracy on average
Figure 5: Few-shot downstream performance on various benchmarks with respect to pretraining iterations for Random and FRAMES.

### v4

# FRAME: Boosting LLMs with A Four-Quadrant Multi-Stage Pretraining Strategy

@@ Abstract
Large language models (LLMs) have significantly advanced human language understanding and generation, with pretraining data quality and organization being crucial to their performance. Multi-stage pretraining is a promising approach, but existing methods often lack quantitative criteria for data partitioning and instead rely on intuitive heuristics. In this paper, we propose the novel Four-quadRAnt Multi-stage prEtraining strategy (FRAME), guided by the established principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. This principle is grounded in two key findings: first, training on high Perplexity (PPL) data followed by low PPL data, and second, training on low PPL difference (PD) data followed by high PD data, both causing the loss to drop significantly twice and performance enhancements. By partitioning data into four quadrants and strategically organizing them, FRAME achieves a remarkable 16.8% average improvement over random across MMLU and CMMLU for the 3B model, effectively boosting LLM performance.
FRAME: Boosting LLMs with
@@ 1 Introduction
(c) Accuracy on CEVAL.
(d) Accuracy on average.
Figure 1: Few-shot downstream performance on various benchmarks with respect to pretraining iterations, for 3B models trained on 1T tokens. The average performance is based on 8 downstream tasks. FRAME achieves 15.3% improvement over Random across MMLU, and 18.2% on CMMLU.
In this paper, we propose a novel Four-quadRAnt Multi-stage prEtraining strategy (FRAME) to boost LLMs’ performance. This strategy is systematically guided by the principle of organizing the pretraining process into four stages to achieve significant loss drops four times. The principle is derived from two key findings based on quantifiable data metrics.
@@
Based on these two major findings, we establish the principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. Specifically, we partition the data into four quadrants based on PPL and PD: Quadrant 1 ($Q_{1}$) with low PD and low PPL, Quadrant 2 ($Q_{2}$) with high PD and low PPL, Quadrant 3 ($Q_{3}$) with low PD and high PPL, and Quadrant 4 ($Q_{4}$) with high PD and high PPL. Through further analysis and experiments, we determine that the optimal strategy is to reorganize the pretraining data in the $Q3\to Q4\to Q1\to Q2$ sequence. To ensure smooth data transitions between stages, we also implement a smoothing process for gradual stage transitions. All data is processed offline to ensure the continuity of model training is not disrupted. Practical evidence shows that FRAME results in loss dropping significantly four times, enhances the models’ emergent abilities, and boosts their performance. As shown in Figure 1, FRAME achieves a significant average improvement of 15.3% on MMLU and 18.2% on CMMLU over random selection, respectively.
Figure 2: The overall framework of FRAME.
In summary, our contributions are as follows: (1) We identify two key findings: first, training on high PPL data followed by low PPL data, and second, training on low PD data followed by high PD data, both causing two significant loss reductions and performance enhancements. (2) Guided by the principle of continuous loss reduction, we propose FRAME, a novel strategy that partitions data into four quadrants and organizes the pretraining process into four stages, resulting in loss dropping significantly four times and boosting model performance. (3) Experiments on 3B model, trained on 1T tokens, demonstrate a 16.8% improvement over uniform sampling in average performance across MMLU and CMMLU.
@@ 2 Four Quadrant Multi-stage Pretraining Strategy
In this section, we first evaluate the effectiveness of PPL and PD as metrics for data organization in two-stage training (Sections 2.1 and 2.2). Based on these insights, we propose FRAME with both metrics(Section 2.3), as shown in Figure 2.
@@ 2.1 Two-stage pretraining guided by PPL
(a) PPL training loss.
(b) PD training loss.
@@
Random 24.8 25.6
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 26.0 26.0
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.6 42.6
@@ 2.2 Two-stage pertaining guided by PD
where $PPL_{M_{w}}(x)$ and $PPL_{M_{s}}(x)$ are the perplexity of the sample $x$ calculated using $M_{w}$ and $M_{s}$, respectively. A low PD indicates similar learning efficiency for both models, while a high PD suggests the sample is more challenging for the weak model.
@@
The distributions of PPL and PD on different domains are analyzed in Appendix B.5, PD exhibits a relatively consistent distribution across varied domains, following a normal distribution with an approximate mean value of 0.3. Compared to PPL, PD offers a better advantage in maintaining data diversity throughout each stage of training.
@@
We train a 100M parameter RM and then calculate the PD for each sample using both the 100M and 1.3B RMs, referred to as PD(100M-1.3B). Using the median value of PD across all data, we divide the training dataset into two subsets with equal token counts: $\text{A}_{\text{PD}}^{\text{low}}$ and $\text{A}_{\text{PD}}^{\text{high}}$. Based on PDPC’s finding that a low-to-high PD ordering achieves better results, we conduct a two-stage training process with $\text{A}_{\text{PD}}^{\text{low}}$ first, followed by $\text{A}_{\text{PD}}^{\text{high}}$, and compare the results to the random setting. We use the evaluation methods described in Section 2.1.
@@
Random 24.8 25.6
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.9 27.2
@@ Results
Figure 3(b) shows the training loss changes under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setup. The loss initially drops rapidly during the low PD phase, then stabilizes, and decreases further with high PD data, eventually falling below the Random model’s loss. It suggests that the first phase sets a beneficial optimization path for the second, helping avoid local optima. Table 2 shows the model’s accuracy under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. Notably, it exceeds the Random setting by 2.1% on MMLU and 1.6% on CMMLU, which validates PD as an effective metric. Further, this leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
@@ 2.3 Four-Quadrant Guided Training Strategy
Inspired by the two key findings about PPL and PD, we establish the principle of organizing pretraining data to achieve significant reductions in training loss. Based on this principle, we introduce a novel Four quadRAnt Multi-stage prEtraining strategy (FRAME), which uses PPL and PD to partition data and reorganize the data sequence to make the training loss drop significantly. The core process of FRAME is shown in Algorithm 1.
Specifically, we train two RMs on the target training set $D$, with the strong model $M_{s}$ having more parameters than the weak model $M_{w}$. Both models are trained on data from the same distribution and under identical settings. We compute PPL using $M_{s}$ and PD using both $M_{s}$ and $M_{w}$ for each data point in $D$. We use the two metrics to partition $\mathcal{D}$ into four quadrants, as shown in Figure 4(a). Our main goal is to ensure that the token numbers of the four quadrants are roughly equal. Therefore, we first determine the PPL threshold to divide the entire dataset into two parts with equal token numbers. Then, for both the high PPL and low PPL parts, we separately find each of their respective PD thresholds to further divide them into two sub-parts with equal token numbers, resulting in four quadrants. During the experiments, there are only slight differences in the PD thresholds for the two PPL subsets, and these dierences did not affect the final experimental results.
The data from four quadrants can be described as follows: Quadrant $\mathbf{Q_{1}}$ contains data that both $M_{s}$ and $M_{w}$ can fit well; Quadrant $\mathbf{Q_{2}}$ contains data that $M_{s}$ learns well, but $M_{w}$ struggles with. Data from this subset is challenging and requires a higher model capacity to understand them; Quadrant $\mathbf{Q_{3}}$ contains data poorly learned by both $M_{s}$ and $M_{w}$; Quadrant $\mathbf{Q_{4}}$ contains data poorly learned by $M_{s}$ and even worse by $M_{w}$.
@@
Figure 4: Four-quadrant partitioning and pretraining losses of four quadrants.
@@
To efficiently organize data from the four quadrants while adhering to constraints, we use a double-loop approach: first, dividing the training into two main phases based on a specific constraint, then further splitting each main phase into two sub-phases according to another constraint. The approach yields two distinct four-stage training strategies: the first strategy follows the sequence $Q_{3}\rightarrow Q_{4}\rightarrow Q_{1}\rightarrow Q_{2}$, while the second adopts $Q_{3}\rightarrow Q_{1}\rightarrow Q_{4}\rightarrow Q_{2}$. Both strategies break the constraints between the second and third stages. However, experiments in Sections 2.1 and 2.2 show that the loss change from low to high PPL is much greater than from high to low PD. Thus, we prioritize the constraints in Equation (2), making the first four-stage training strategy $\mathbf{Q_{3}\to Q_{4}\to Q_{1}\to Q_{2}}$ the better choice.
Algorithm 1 Ordered dataset construction via FRAME
1:Input: pretraining data $D$, batch size $N$, steepness $a$
2:Output: ordered dataset $S_{\text{frame}}$
3:function MergeDatasets($D_{1}$, $D_{2}$, RandomSample)
4:     Total training steps $m\leftarrow\frac{|D_{1}|+|D_{2}|}{N}$
5:     if RandomSample then
6:         Randomly shuffle $D_{1}$ and $D_{2}$
7:     end if
8:     Initialize $l\leftarrow 0$, $r\leftarrow 0$, $Q_{\text{merged}}\leftarrow[]$
9:     for i = 1 todo
10:         Calculate completion ratio $p\leftarrow\frac{i}{m}$
11:         Calculate proportion $f(p)\leftarrow\frac{1}{1+\exp(a(p-0.5))}$
12:         $B_{1}\leftarrow D_{1}[l:l+f(p)N]$
13:         $B_{2}\leftarrow D_{2}[r:r+(1-f(p))N]$
14:         Batch $B_{i}=B_{1}\cup B_{2}$
15:         $l\leftarrow l+f(p)N$, $r\leftarrow r+(1-f(p))N$
16:         Add all the samples in $B_{i}$ to $Q_{\text{merged}}$
17:     end for
18:     return $Q_{\text{merged}}$
19:end function
20:Train RMs on i.i.d. subset of $D$
21:Calculate PPL and PD for all samples in $D$ using RMs
22:Divide into 4 quadrants: $Q_{1}$, $Q_{2}$, $Q_{3}$, $Q_{4}$, based on PPL and PD thresholds.
23:$S_{34}\leftarrow\textsc{MergeDatasets}(Q_{3},Q_{4},\text{True})$
24:$S_{12}\leftarrow\textsc{MergeDatasets}(Q_{1},Q_{2},\text{True})$
25:$S_{\text{frame}}\leftarrow\textsc{MergeDatasets}(S_{34},S_{12},\text{False})$
Formulation of FRAME with Stage Transition
As illustrated in Figure 6(b), direct stage transitions cause performance fluctuations. We aim to facilitate a smooth transition between stages.
We start by outlining the smoothing process $f_{merge}$ for two-stage training: given the need to train on mixed data from two subsets $D_{1}$ and $D_{2}$, let $i$ be the current training step and $m$ the total number of training steps. The completion ratio is defined as $p=\frac{i}{m}$. The sampled batch $B_{i}$ should satisfy the following condition:
$B_{i}=\{x\mid x\sim D_{1}\}_{f(p)\cdot N}\cup\{x\mid x\sim D_{2}\}_{(1-f(p))\cdot N}$ (4)
where $N$ represents the batch size, and $f(p)$ denotes the proportion of samples from $D_{1}$. Inspired by PDPC Zhang et al. (2025), we employ the S-shape function as $f(p)$:
$f(p)=\frac{1}{1+\exp(a(p-0.5))}$ (5)
@@
where $a$ controls the steepness of the curve. Unlike PDPC, we merely utilize the S-shape function for smoothing during stage transitions. Therefore, we use a larger $a=35$ instead of $a=10$ as in PDPC, to achieve a steeper function curve, as illustrated in Appendix B.4.
In the four-stage training of FRAME, we initially obtain $S_{34}=f_{merge}(Q_{3},Q_{4})$ and $S_{12}=f_{merge}(Q_{1},Q_{2})$. During this phase, samples are randomly selected from the two data sources for each batch. Subsequently, we construct $S_{frame}=f_{merge}(S_{34},S_{12})$. In this phase, samples must be drawn in the original sequence of $S_{12}$ and $S_{34}$ to maintain the order of $Q_{3}\rightarrow Q_{4}\rightarrow Q_{1}\rightarrow Q_{2}$.
@@
Notably, FRAME only organizes the given data without performing selection, which can serve as the final data preprocessing step before pretraining, without disrupting the continuity of training.
@@ 3 Experiments
Method Metric Order ARC-E ARC-C SciQ HellaSw. PIQA AVG.
- - Random 56.5 23.6 85.8 34.2 67.3 53.5
Sequential PD High2Low 54.7$\downarrow$1.8 21.8$\downarrow$1.8 87.1$\uparrow$1.3 33.7$\downarrow$0.5 67.8$\uparrow$0.5 53.0$\downarrow$0.5
PD Low2High 56.1$\downarrow$0.4 21.3$\downarrow$2.3 86.2$\uparrow$0.4 34.4$\uparrow$0.2 67.6$\downarrow$0.2 53.1$\downarrow$0.4
PPL High2Low 45.5$\downarrow$11.0 20.6$\downarrow$3.0 71.2$\downarrow$14.6 30.3$\downarrow$3.9 63.7$\downarrow$3.6 46.3$\downarrow$7.2
PPL Low2High 47.8$\downarrow$8.7 17.9$\downarrow$5.7 72.7$\downarrow$13.1 29.1$\downarrow$5.1 62.4$\downarrow$4.9 46.0$\downarrow$7.5
Qu.Edu High2Low 57.2$\uparrow$0.7 26.4$\uparrow$2.8 85.4$\downarrow$0.4 33.0$\downarrow$1.2 66.2$\downarrow$1.1 53.6$\uparrow$0.1
Qu.Edu Low2High 56.8$\uparrow$0.3 26.0$\uparrow$2.4 84.1$\downarrow$1.7 33.5$\downarrow$0.7 67.9$\uparrow$0.6 53.7$\uparrow$0.2
Preference CL PPL S.R. 56.1$\downarrow$0.4 24.1$\uparrow$0.5 87.8$\uparrow$2.0 33.9$\downarrow$0.3 67.4$\uparrow$0.1 53.9$\uparrow$0.4
@@
Qu.Edu S.R 56.7$\uparrow$0.2 24.9$\uparrow$1.3 86.2$\uparrow$0.4 33.6$\downarrow$0.6 66.9$\downarrow$0.4 53.7$\uparrow$0.2
Qu.Edu S. 55.5$\downarrow$1.0 24.8$\uparrow$1.2 87.8$\uparrow$2.0 34.0$\uparrow$0.2 67.4$\uparrow$0.1 53.9$\uparrow$0.4
PD S.R. 56.7$\uparrow$0.2 24.9$\uparrow$1.3 86.2$\uparrow$0.4 33.6$\downarrow$0.6 67.4$\uparrow$0.1 53.8$\uparrow$0.3
PDPC PD S. 57.3$\uparrow$0.8 26.6$\uparrow$3.0 87.9$\uparrow$2.1 33.7$\downarrow$0.5 68.0$\uparrow$0.7 54.7$\uparrow$1.2
FRAME PPL&PD - 62.9$\uparrow$6.4 26.5$\uparrow$2.9 90.5$\uparrow$4.7 38.4$\uparrow$4.2 69.6$\uparrow$2.3 57.6$\uparrow$4.1
Table 3: Downstream tasks results for different settings on 1.3B models. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: AVG. = Average, S.=S-shape Function, S.R.=S-shape Reverse Function.
Method MMLU CMMLU CEVAL BBH ARC-E ARC-C HellaSw. PIQA AVG.
Random 27.7 27.5 27.2 27.9 68.6 33.7 49.4 76.0 42.3
PDPC 35.8$\uparrow$8.1 35.6$\uparrow$8.1 36.1$\uparrow$8.9 25.7$\downarrow$2.2 69.7$\uparrow$1.1 35.8$\uparrow$2.1 49.9$\uparrow$0.5 76.3$\uparrow$0.3 45.6$\uparrow$3.3
$Q_{3}\to Q_{1}\to Q_{4}\dashrightarrow Q_{2}$ 25.8$\downarrow$1.9 26.4$\downarrow$1.1 25.5$\downarrow$1.6 28.9$\uparrow$1.0 68.3$\downarrow$0.3 35.2$\uparrow$1.5 48.3$\downarrow$1.1 75.4$\downarrow$0.6 41.7$\downarrow$0.6
FRAME 43.0$\uparrow$15.3 45.7$\uparrow$18.2 44.0$\uparrow$16.8 27.9   0.0 71.0$\uparrow$2.4 36.5$\uparrow$2.8 50.2$\uparrow$0.8 76.9$\uparrow$0.9 49.4$\uparrow$6.1